OpenStudy (anonymous):

find the equation of the line containing the given pair of points. (-5,0) & (0,4)

OpenStudy (oaktree):

Find the slope by\[(y _{2}+y _{1})\div(x _{2}-x _{1})\] where the y's and x's are the ones from your coordinates.

OpenStudy (oaktree):

Then plug in a set of x and y into y=mx+b (where m=slope and b=y-intercept) and also plug in the slope to solve for b. Then once you have b write y=mx+b but write the actual m and b, but leave the x and y as x and y.

OpenStudy (oaktree):

So your slope is 4/5

OpenStudy (oaktree):

Use (0,4) and plug in to get b=16/5

OpenStudy (oaktree):

So your equation is y=4/5x+16/5

OpenStudy (anonymous):

thank you